.52h of a cheese puff is burned under an aluminum can filled with 50.1 ML of water. The water temperature in the can increases by 2.3°C. how many calories did the cheese puff release?
1
Expert's answer
2017-03-27T16:08:06-0400
In assumption that the energy released is not dissipated and all transferred in the energy of the water heating, we can write the following equation:

Q_combustion=Q_water

The change of the temperature of water is related to the energy of heating as:

Q_water=cm∆T,

where c is the specific heat of water: c=1 cal kg^(-1) ℃^(-1). Assuming 1g/ml density of water, the mass of the water is 50.1 ml* 1 g/ml = 50.1 g. Then, the energy released is:

Q_combustion=Q_water=1(cal kg^(-1) ℃^(-1) )·50.1(g)·10^(-3) (kg g^(-1))·2.3(℃)= 0.115 (cal)

Answer: The cheese puff released 0.115 cal.